Diffusion Morning Report 02/20/09

Ted Berg tberg at stanford.edu
Fri Feb 20 06:23:15 PST 2009


Good Morning All,
    Looking good this morning Still checking out SiGe. TEOS 2 pump is 
dead and out for repair. Ebam still struggling. No other issues at this 
time.

  	  	  	  	  	 
  	Date: 	2/20/09 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	down 	 clean Completed Pump dead 	0/30 	  	 
Tylan 8-SiGe 	yellow 	Hardware check temp cal 	1.12/30 	  	 
Tylan 9-Poly 	up 	  	.13/30 	  	 
Tylan 10-Nitride 	up 	  	1.59/30 	  	 
Tylan 11-LTO 	up 	  	25.63/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	up 	  	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	  	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	up 	  	   	  	 
Ebeam 	down 	tip changed looking for beam 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	B on line 47 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20090220/e1cc3d22/attachment.html>


More information about the diffusion mailing list